OpenStudy (anonymous):

10 is multiplied by the sum of 3 and 2. The product is divided by 2/and then 5 is added @Mertsj @DarthTony @

OpenStudy (anonymous):

10 is multiplied by the sum of 3 and 2. Let's just call this answer "x" \[x=10*(3+2)=50\] The product (x, which is 50) is divided by 2 and then 5 is added: \[(x/2)+5=(50/2)+5=25+5=30\]

OpenStudy (anonymous):

all i need is u to write out a problem not anwser

OpenStudy (mertsj):

Let's break it down: 10 is multiplied by the sum of 3 and 2: 10(3+2) The product (remember the answer to a multiplication problem is a product) so the product is 10(3+2) It is divided by 2 so we have: \[\frac{10(3+2)}{2}\]

OpenStudy (mertsj):

And then 5 is added so we have: \[\frac{10(3+2)}{2}+5\]

OpenStudy (anonymous):

\[\frac{ 10(3+2) }{ 2 }+5\]

OpenStudy (anonymous):

thats what i thought it was u wanna help me with more